OpenStudy (anonymous):

A bag contains a total of 14 batteries, of which five are defective. Selecting two at random, without replacement, determine the probability that none of the batteries you select are good. 1/2 1/7 10/91 5/14

OpenStudy (anonymous):

Looks like it's the 3rd one: (5/14)*(4/13) = 10/91

OpenStudy (anonymous):

There are 5 defective ones to choose at first. So assuming you pick a bad one, there are 4 defective ones left out of a total of 13. The chance of both happening is the product of the individuals: \[\frac{5}{14}*\frac{4}{13} = \frac{10}{91}\]